βˆ‘ (2 𝑗+1 βˆ’ 2 𝑗 ) 8 𝑗=0Β 


1
Expert's answer
2021-12-07T11:40:00-0500

"\displaystyle\sum_{j=0}^8(2 𝑗+1 − 2 𝑗 )=\displaystyle\sum_{j=0}^81=9"